# Communication Constellation Design of Minimum Number of Satellites with Continuous Coverage and Inter-Satellite Link

### v1

@@ Abstract
Recently, the advancement of research on distributed space systems that operate a large number of satellites as a single system urges the need for the investigation of satellite constellations. The communication constellation can be used to construct global or regional communication networks using inter-satellite and ground-to-satellite links. This paper studies the two challenges of the communication constellation, continuous coverage and inter-satellite link connectivity. The bounded Voronoi diagram and APC decomposition are presented as continuous coverage analysis methods. For the continuity analysis of the inter-satellite link, the relative motion between the adjacent orbital planes derives the analytic solutions. The Walker-Delta constellation and common ground-track constellation design methods are introduced as examples to verify the analysis methods. The common ground-track constellation is classified into a quasi-symmetric and optimal constellation. The optimal common ground-track constellation is optimized by the BILP algorithm. As a result, the simulation results compare the performance of the communication constellations according to various design methods.
@@ 1 Introduction
In recent years, the feasibility of low Earth orbit communication satellites has attracted attention due to the trend of small satellites. The examples include Space-X’s Starlink, Eutelsat’s OneWeb, and Amazon’s Kuiper project[1]. Their orbits are designed to provide communication links across the whole Earth. A satellite constellation operates several satellites with distinct orbits as a single system. Compared to conventional geostationary satellites, low Earth orbit satellites have a smaller coverage size and shorter orbital periods, which results in degraded spatial and temporal coverage performance. Therefore, the low Earth orbit constellations operate several satellites to overcome this limitation while minimizing the number of satellites. If the coverage is continuous for the area of interest, the communication constellation can continuously provide communication links. In addition, when the inter-satellite links are connected, the link is not disconnected even by the orbital motion of the satellite and the Earth’s rotation effect. In this paper, the low Earth orbit communication constellation design problem is interpreted as the problem of achieving continuous coverage and inter-satellite links.
The sections are presented in order of constellation design methods, coverage analysis methods, relative motion in adjacent orbital planes, simulation results, and conclusion. The constellation design methods describe the basis of Walker and common ground-track constellations. The Walker constellation first introduces the concept of seed satellite and design parameters and defines the orbital elements. The pattern repetition period and duplicate allocation of orbital elements are explained as two important characteristics of the Walker-Delta constellation and expressed by the design parameters. The repeat ground-track orbit is the seed satellite’s orbit for the common ground-track constellation. The quasi-symmetric method configures the satellites so that the spacings between the adjacent satellites are almost equal. The BILP method optimizes the satellite’s configuration to satisfy the coverage requirement while minimizing the total number of satellites. The coverage analysis methods section first explains a useful concept, the geometry of Earth coverage. Then, Voronoi tessellation and APC decomposition are described with the references. The relative motion in adjacent orbital planes derives key formulae to analyze the bounds of the relative distance. The simulation results present the continuous coverage analysis for Walker-Delta, quasi-symmetric, and BILP constellations and inter-satellite link continuity analysis for three constellations. The last section summarizes and concludes the contents of this paper.
@@ 2.1 Walker Constellation
The Walker constellation is the geometric design method that configures the orbits symmetrically[2, 3]. The seed satellite is the one that determines the common orbital elements of the constellation. For the Walker constellation, the satellites are configured in circular orbits with the equal semi-major axis ($a$), inclination ($i$), and argument of perigee ($\omega$) of the seed satellites. The two types of Walker constellations are classified depending on the inclination. The Walker-Star constellation is designed based on the polar orbit and achieves global coverage. The Walker-Delta constellation, on the other hand, has an inclination under 90 degrees and covers the mid-latitude and equator region. Therefore, this classification is only relevant to the latitude of the area of interest (AoI).
Figure 1: Geometric description of Walker-Delta constellation orbital elements
The three design parameters are the total number of satellites ($T$), the number of orbital planes ($P$), and the phasing parameter ($F\in\{0,1,...,P-1\}$). The number of satellites per orbital plane ($S=T/P$) is an auxiliary parameter to prevent confusion. The right ascension of ascending node (RAAN, $\Omega_{m}$) and the mean anomaly ($M_{m,n}$) of the $n$th satellite in the $m$th orbital plane ($SAT_{m,n}$) are defined as Eqs. (1) and (2) and depicted in Figure 1.
@@
From the equation (1), the RAAN is equally spaced by $\Delta\Omega=360/P\deg$ in the range of $\Omega_{m}\in\left[0,360\right]\deg$. The intraplane spacing $\Delta M=360/S\deg$ generates the symmetric location in the orbital plane in the range of $M\in\left[0,360\right]\deg$ and is equivalent to the relative angular distance between the satellites in the orbital plane. The relative argument of latitude ($\Delta u$) in the adjacent orbital planes is derived as $\Delta u=360/T\cdot F\deg$ from Eq. (2).
@@ 2.1.1 Pattern Repetition Period
The Walker-Delta pattern has the geometric characteristics that the intra- and inter-plane angular spacings are uniform. This geometric symmetricity determines the pattern repetition period. The investigation of the patterns in the geographical coordinate system enhances understanding of the pattern repetition period. The reference [4] derived the formulae of pattern repetition period. The pattern unit (PU) more briefly describes the orbital elements of Walker-Delta constellation. The definition of the pattern unit is as below[2, 3]:
@@
The RAAN and mean anomaly in Eqs. (1) and (2) are reorganized in pattern unit as
@@
The time interval that the mean anomaly increases for $F\text{PU}$ is the first pattern repetition period $t_{F}$ and defined as
@@
where $\omega_{orb}$ is the orbital angular speed. Assuming the twobody motion, the RAAN and mean anomaly at $t_{F}$ are derived as
@@
where $t_{0}$ is the epoch time.
From the Eqs. (7) and (8), the mean anomaly of the $n$th satellite in $m$th orbital plane at $t_{F}$ $M_{m,n}\left(t_{F}\right)$ is the same with the one of the $n$th satellite in $m+1$th orbital plane at $t_{0}$ $M_{m+1,n}\left(t_{0}\right)$. Thus, the constellation pattern observed at $t_{F}$ appears as a pattern shifted $\Delta\Omega\deg$ westward at time $t_{0}$.
@@
On the other hand, the mean anomaly propagates $P\text{PU}$ during $t_{P}$ and it defines the second pattern repetition period as
@@
The RAAN and mean anomaly at $t_{P}$ are derived as
@@
Let us define the set of satellites in the $m$th orbital plane as $\textbf{N}=\left\{n\mid n=1,2,...,S\right\}$. Then, the equations (10) and (11) derives
@@
Therefore, the constellation pattern at $t_{P}$ appears exactly the same as the pattern at $t_{0}$.
@@ 2.1.2 Duplicate Allocation of Orbital Elements
The Walker-Delta design method allocates the six unique orbital elements to each satellite corresponding to the orbital state of six position and velocity elements. It suggests the possibility of the duplicate positions of the satellites and implies the redundant number of satellites or the collisions between satellites. The conditions of duplicate orbital elements of Walker-Delta constellation is investigated in the reference 4 as
@@
The equation (13) can be expressed by Walker-Delta design parameters as
@@
In the design procedure, the Walker-Delta design parameters that accommodate the Eq. (14) should be avoided.
@@ 2.2.1 Repeat Ground-track Orbit
The repeat ground-track orbit (RGT) is defined as the orbit that traces the same ground-track within a specified time interval. The two main parameters that determines the RGT orbit are the revolutions to repeat ($N_{P}$) and the days to repeat ($N_{D}$)[5]. For example, if the revolutions to repeat is $14$ and the days to repeat is $1$, the ground-track crosses (ascends or descends) the equator 14 times during 1 day. From this, the period ratio ($\nu$), the RGT design parameter, is formulated as
@@
The period ratio can also be described by the satellite nodal period ($T_{S}$) and the nodal period of greenwich ($T_{G}$). The change of the orbital elements by $J_{2}$ effect induces the changes in $T_{S}$ and $T_{G}$ as
@@
where $\omega$ is the argument of the perigee, $\Omega$ is the right ascension of the ascending node (RAAN), $M$ is the mean anomaly, $\omega_{E}$ is the Earth’s rotation speed.
The orbital elements are formulated as Eqs. (18), (19), and (20)
@@
From the definition of the period ratio, the introduction of Eqs. (16) and (17) derives the Eq. (15) with respect to the orbital elements.
@@
From the Eqs. (18), (19), and (20), the orbital elements have their arguments as $\dot{\omega}=\dot{\omega}\left(a,i,e\right)$, $\dot{M}=\dot{M}\left(a,i,e\right)$, and $\dot{\Omega}=\dot{\Omega}\left(a,i,e\right)$. It organizes the arguments of $T_{S}$, $T_{G}$, and $\nu$ as
@@
From this, the RGT orbital design algorithm can be derived. Given the specific $\nu$, inclination ($i$), and eccentricity ($e$), the algorithm calculates the corresponding semi-major axis (a). The equation (15) implies that $\nu$ determines the number of revolutions per a certain period and suggests that $\nu$ is related to the orbital period and the semi-major axis. As a result, when $e$ and $i$ are specified, one unique semi-major axis is derived from a $\nu$ and vice versa. For example, if the set of RGt orbital elements is given as $\left(\nu,i,e\right)=\left(14,42\deg,0\right)$, then $a$ is $7201.90km$. The RGT orbital elements can be described in a different way, such as $\left(a,i,e\right)=\left(7201.90km,42\deg,0\right)$, then $\nu$ is uniquely determined as $14$.
@@ 2.2.2 Common Ground-track Constellation
The RGT orbit designed following the Eq. (21) has a set of $\left(\nu,i,e\right)$ or $\left(a,i,e\right)$ with an arbitrary set of $\left(\omega,\Omega,M\right)$. From here on, the orbital elements of RGT orbit is described as $\left(\nu,i,e\right)$ except a specific purpose. This implies that a numerous number of satellites could trace the same ground-track and introduces the concept of common ground-track (CGT) constellation. The CGT constellation is designed following the the three procedures below[6, 7]:
@@
(2) Choose an arbitrary $\omega$ so that all the satellites have the same $\left(\nu,i,e,\omega\right)$.
(3) Given the CGT constellation’s total number of satellite ($T$), the $k$th satellite’s RAAN and mean anomaly $\left(\Omega_{k},M_{k}\right)$ satisfies Eq. (25)
@@
The CGT constellation design problem is concluded as the configuration method of $\left(\Omega_{k},M_{k}\right)$ following the procedure (3). The reference  [7] introduces two methods: the quasi-symmetric and binary integer programming (BILP) method.
@@ 2.2.3 Quasi-symmetric Method
The simulation time ($T_{sim}$) is discretized by the step size $t_{step}$ and is assumed to be the integer multiple of the repetition period. Then, it can be formulated as $T_{sim}=L\cdot t_{step}$. The continuous time variable $t\in\left[0,T_{sim}\right]$ is converted into the discretized time variable $\tau\in\left\{0,1,...,L-1\right\}$. The configuration of the satellites of the CGT constellation can be expressed as the time shifted seed satellites because all the satellites have the same ground-track traces. Therefore, the constellation pattern vector $x\left[\tau\right]$ can be defined as
@@
In the discretized time domain, the constellation pattern vector has the length $L$. If the total number of the constellation is $T$, the spacing constant $\xi$ is defined as
$\xi\triangleq\frac{L}{T}$ (27)
If $\xi$ is an integer, $\tau_{k}$ is spaced equally and the satellites are configured symmetrically. On the other hand, if $T$ is not a divisor of $L$, then $\xi$ is not an integer that makes the index $\tau_{k}$ also a rational number. In this case, only a quasi-symmetric configuration is possible. The formulation for both symmetric and quasi-symmetric constellation pattern vector $\overline{x}\left[\tau\right]$ is defined as
@@
where $nint$ is the nearest integer function.
@@ 2.2.4 Binary Integer Linear Programming Method
Because the domain of the constellation pattern vector is defined as binary, BILP method is introduced as the optimization algorithm. The BILP algorithm is one of the variants of the linear programming method that optimizes the linearized objective function with the constraints and boundaries. The problem statement of the linear programming can be generalized as[8]
@@
In addition to Eq. (29), the domain of the decision variable x needs to be defined. The domain sets of $\mathbb{R}_{\geq 0}$, $\mathbb{Z}_{\geq 0}$, and $\mathbb{Z}_{2}$ induce the pure real, integer, and binary integer linear programming. The mixed domains are also possible that mixed integer linear programming have the real and integer decision variables. Thus, the BILP has the same problem statement but constraints the domain of decision variables as
@@
The objective of the optimization is to obtain the constellation pattern vector that minimizes the number of satellites while satisfying the coverage requirement. Since the summation of $x\left[\tau\right]$ is equal to $T$ by the definition of the constellation pattern vector in Eq. (26), the problem is formulated as
@@
where $j$ is the index for the $j$th grid, $\mathcal{J}$ is the set of grids in the area of interest and $\mathbb{Z}_{2}$ is the binary integer number set. The matrix $\textbf{V}_{0,j}\in\mathbb{Z}_{2}^{L\times L}$ is a seed-satellite access profile circulant matrix and will be addressed in detail in the next chapter.

### v2

@@ Abstract
The recent advancement in research on distributed space systems that operate a large number of satellites as a single system urges the need for the investigation of satellite constellations. Communication constellations can be used to construct global or regional communication networks using inter-satellite and ground-to-satellite links. This study examines two challenges of communication constellations: continuous coverage and inter-satellite link connectivity. The bounded Voronoi diagram and APC decomposition are presented as continuous coverage analysis methods. For continuity analysis of the inter-satellite link, the relative motion between adjacent orbital planes is used to derive analytic solutions. The Walker–Delta constellation and common ground-track constellation design methods are introduced as examples to verify the analysis methods. The common ground-track constellations are classified into quasi-symmetric and optimal constellations. The optimal common ground-track constellation is optimized using the BILP algorithm. The simulation results compare the performance of the communication constellations according to various design methods.
@@ 1 Introduction
In recent years, the feasibility of low Earth orbit communication satellites has attracted attention owing to the trend of small satellites. Examples include Space-X’s Starlink, Eutelsat’s OneWeb, and Amazon’s Kuiper projects [1]. Their orbits are designed to provide communication links across the Earth. A satellite constellation operates several satellites with distinct orbits in a single system. Compared with conventional geostationary satellites, low Earth orbit satellites have a smaller coverage size and shorter orbital periods, resulting in degraded spatial and temporal coverage performance. Therefore, low Earth orbit constellations operate several satellites to overcome this limitation while minimizing the number of satellites. If the coverage is continuous in the area of interest, the communication constellation can continuously provide communication links. In addition, when the inter-satellite links are connected, the link is not disconnected even by the orbital motion of the satellite and the Earth’s rotation effect. In this study, the low Earth orbit communication constellation design problem was interpreted as the problem of achieving continuous coverage and inter-satellite links.
The sections are presented in the following order: constellation design methods, coverage analysis methods, relative motion in adjacent orbital planes, simulation results, and conclusions. Constellation design methods describe the basis of Walker and common ground-track constellations. The Walker constellation first introduces the concept of a seed satellite and its design parameters, and defines the orbital elements. The pattern repetition period and duplicate allocation of orbital elements are explained as two important characteristics of the Walker-Delta constellation, and are expressed by the design parameters. The repeat ground-track orbit is the seed satellite orbit for a common ground-track constellation. The quasi-symmetric method configures satellites such that the spacing between adjacent satellites is almost equal. The BILP method optimizes the satellite configuration to satisfy the coverage requirements while minimizing the total number of satellites. The coverage analysis methods section first explains a useful concept: the geometry of the Earth’s coverage. Then, Voronoi tessellation and APC decomposition are described with the references. The relative motion in adjacent orbital planes derives the key formulae to analyze the bounds of the relative distance. The simulation results present the continuous coverage analyses for the Walker-Delta, quasi-symmetric, and BILP constellations, and the inter-satellite link continuity analyses for the three constellations. The last section summarizes and concludes the contents of this paper.
@@ 2.1 Walker Constellation
The Walker constellation is a geometric design method that symmetrically configures the orbits [2, 3]. The seed satellite determines the common orbital elements of the constellation. For the Walker constellation, the satellites are configured in circular orbits with the same semi-major axis ($a$), inclination ($i$), and argument of perigee ($\omega$) of the seed satellites. The two types of Walker constellations are classified according to their inclinations. The Walker-Star constellation is designed based on a polar orbit and achieves global coverage. On the other hand, the Walker-Delta constellation has an inclination under 90 degrees and covers the mid-latitude and equator regions. Therefore, this classification is relevant only to the latitude of the area of interest (AoI).
Figure 1: Geometric description of the orbital elements of the Walker-Delta constellation
The three design parameters are the total number of satellites ($T$), number of orbital planes ($P$), and phasing parameter ($F\in\{0,1,...,P-1\}$). The number of satellites per orbital plane ($S=T/P$) is an auxiliary parameter used to prevent confusion. The right ascension of ascending node (RAAN, $\Omega_{m}$) and mean anomaly ($M_{m,n}$) of the $n$th satellite in the $m$th orbital plane ($SAT_{m,n}$) are defined by Eqs. (1) and (2), as shown in Figure 1.
@@
From the equation (1), the RAAN is equally spaced by $\Delta\Omega=360/P\deg$ in the range $\Omega_{m}\in\left[0,360\right]$ $\deg$. The intraplane spacing $\Delta M=360/S\deg$ generates a symmetric location in the orbital plane in the range $M\in\left[0,360\right]\deg$ and is equivalent to the relative angular distance between the satellites in the orbital plane. The relative argument of the latitude ($\Delta u$) in the adjacent orbital planes is derived as $\Delta u=360/T\cdot F\deg$ from Eq. (2).
@@ 2.1.1 Pattern Repetition Period
The Walker-Delta pattern has geometric characteristics such that the intra- and inter-plane angular spacings are homogeneous. This geometric symmetry determines the pattern repetition period. The investigation of patterns in a geographical coordinate system enhances the understanding of the pattern repetition period. Reference  4 derived the formulas for the pattern repetition period. The pattern unit (PU), which is introduced to understand the orbital configuration of the Walker-Delta constellation, is defined as follows [2, 3]:
@@
The RAAN and mean anomaly in Eqs. (1) and (2) are reorganized in PU as
@@
The time interval in which the mean anomaly increases for $F$ PU is the first pattern repetition period $t_{F}$ and defined as
@@
where $\omega_{orb}$ denotes the orbital angular speed. Assuming the twobody motion, the RAAN and mean anomaly at $t_{F}$ are derived as
@@
where $t_{0}$ denotes the epoch time.
From Eqs. (7) and (8), the mean anomaly of the $n$th satellite in $m$th orbital plane at $t_{F}$, $M_{m,n}\left(t_{F}\right)$, is the same as the one of the $n$th satellite in $(m+1)$th orbital plane at $t_{0}$, $M_{m+1,n}\left(t_{0}\right)$. Thus, the constellation pattern observed at $t_{F}$ appears as a pattern at $t_{0}$ that is shifted westward as $\Delta\Omega\deg$.
@@
On the other hand, the mean anomaly propagates $P$ PU during $t_{P}$ and defines the second pattern repetition period as
@@
The RAAN and the mean anomaly at $t_{P}$ are derived as
@@
Let us define the set of satellites in the $m$th orbital plane as $\textbf{N}=\left\{n\mid n=1,2,...,S\right\}$. Then, Equations (10) and (11) derive
@@
Therefore, the constellation pattern at $t_{P}$ appears identical to the one at $t_{0}$.
@@ 2.1.2 Duplicate Allocation of Orbital Elements
The Walker-Delta design method allocates the six unique orbital elements to each satellite, and the six orbital elements correspond to the unique orbital state of the six positions and velocity elements. This suggests the possibility of duplicate satellite positions and implies the collision between satellites. The conditions of the duplicate orbital elements of the Walker-Delta constellation are investigated in reference 4 as
@@
Equation (13) can be expressed by the Walker-Delta design parameters as follows:
@@
Therefore, the Walker-Delta design parameters that accommodate Eq. (14) must be avoided in the design procedure.
@@ 2.2.1 Repeat Ground-track Orbit
The repeat ground-track (RGT) orbit is an orbit that traces the same ground-track within a specific time interval. The two main parameters that determine the RGT orbit are the number of revolutions to repeat ($N_{P}$) and the number of days to repeat ($N_{D}$)[5]. For example, if the number of revolutions to repeat is $14$ and the number of days to repeat is $1$, the ground-track crosses (ascends or descends) the equator 14 times in one day. Thus, the period ratio ($\nu$), which is the RGT design parameter, is formulated as
@@
The period ratio can also be described by the satellite nodal period ($T_{S}$) and the nodal period of Greenwich ($T_{G}$). The change in the orbital elements due to $J_{2}$ effect induces changes in $T_{S}$ and $T_{G}$ as
@@
where $\omega_{E}$ is Earth’s rotation speed.
The orbital elements are formulated using Eqs. (18), (19), and (20):
@@
Based on the definition of period ratio, Eqs. (16) and (17) derive Eq. (15) with respect to the orbital elements as:
@@
From Eqs. (18), (19), and (20), the orbital elements have arguments as $\dot{\omega}=\dot{\omega}\left(a,i,e\right)$, $\dot{M}=\dot{M}\left(a,i,e\right)$, and $\dot{\Omega}=\dot{\Omega}\left(a,i,e\right)$. It organizes the arguments of $T_{S}$, $T_{G}$, and $\nu$ as:
@@
From this, the RGT orbital design algorithm can be derived. Given the specific $\nu$, inclination ($i$), and eccentricity ($e$), the algorithm calculates the corresponding semi-major axis (a). Equation (15) implies that $\nu$ determines the number of revolutions per period and suggests that $\nu$ is related to the orbital period and semi-major axis. Consequently, when $e$ and $i$ are specified, one unique semi-major axis is derived from $\nu$ and vice versa. For example, if the set of RGT orbital elements is given as $\left(\nu,i,e\right)=\left(14,42\deg,0\right)$, then $a$ is $7201.90km$. As a result, the RGT orbital elements can be described in two different ways, such as $\left(a,i,e\right)=\left(7201.90km,42\deg,0\right)$, and then $\nu$ is uniquely determined as $14$.
@@ 2.2.2 Common Ground-track Constellation
The RGT orbit that is designed according to Eqs. (21) contains a set of $\left(\nu,i,e\right)$ or $\left(a,i,e\right)$ with an arbitrary set of $\left(\omega,\Omega,M\right)$. From here on, the orbital elements of the RGT orbit are denoted as $\left(\nu,i,e\right)$ except for specific purposes. This implies that a numerous number of satellites can trace the same ground-track and introduces the concept of a common ground-track (CGT) constellation. The CGT constellation is designed following the three procedures below[6, 7]:
@@
(2) Choose an arbitrary $\omega$ so that all satellites have the same $\left(\nu,i,e,\omega\right)$.
(3) Given the CGT constellation’s total number of satellites ($T$), the $k$th satellite’s RAAN and mean anomaly $\left(\Omega_{k},M_{k}\right)$ satisfy Eq. (25)
@@
The CGT constellation design problem is concluded as the configuration method of $\left(\Omega_{k},M_{k}\right)$ following procedure (3). The reference  7 introduces two methods, quasi-symmetric and binary integer programming (BILP).
@@ 2.2.3 Quasi-symmetric Method
The simulation time ($T_{sim}$) is discretized by the step size $t_{step}$ and is assumed to be an integer multiple of the repetition period. Thus, this can be formulated as $T_{sim}=L\cdot t_{step}$. The continuous time variable $t\in\left[0,T_{sim}\right]$ is converted into the discretized time variable $\tau\in\left\{0,1,...,L-1\right\}$. The configuration of the satellites in the CGT constellation can be expressed as time-shifted seed satellites because all satellites have the same ground-track traces. Therefore, the constellation pattern vector $x\left[\tau\right]$ can be defined as follows:
@@
In the discretized time domain, the constellation pattern vector has length $L$. If the total number of constellations is $T$, then the spacing constant $\xi$ is defined as follows:
$\xi\triangleq\frac{L}{T}.$ (27)
If $\xi$ is an integer, $\tau_{k}$ is equally spaced and the satellites are configured symmetrically. In contrast, if $T$ is not a divisor of $L$, then $\xi$ is not an integer that makes the index $\tau_{k}$ a rational number. In this case, only a quasi-symmetric configuration is possible. The formulation for both symmetric and quasi-symmetric constellation pattern vectors $\overline{x}\left[\tau\right]$ is defined as
@@
where $nint$ denotes the nearest integer function.
@@ 2.2.4 Binary Integer Linear Programming Method
Because the domain of the constellation pattern vector is defined as binary, the BILP method is introduced as an optimization algorithm. The BILP algorithm is a variant of the linear programming method that optimizes a linearized objective function with constraints and boundaries. The problem statement of the linear programming can be generalized as[8]
@@
In addition to Eq. (29), the domain of the decision variable x must be defined. The domain sets $\mathbb{R}_{\geq 0}$, $\mathbb{Z}_{\geq 0}$, and $\mathbb{Z}_{2}$ induce the pure real and binary integer linear programming. It is also possible for mixed-integer linear programming to include both real and integer decision variables. Thus, the BILP has the same problem statement but constrains the domain of the decision variables as
@@
The objective of optimization is to obtain a constellation pattern vector that minimizes the number of satellites while satisfying the coverage requirement. Because the summation of $x\left[\tau\right]$ is equal to $T$ by the definition of the constellation pattern vector in Eq. (26), the problem is formulated as
@@
where $j$ is the index for the $j$th grid, $\mathcal{J}$ is the set of grids in the area of interest, and $\mathbb{Z}_{2}$ is the binary integer number set. The matrix $\textbf{V}_{0,j}\in\mathbb{Z}_{2}^{L\times L}$ is a seed-satellite access profile circulant matrix, which is addressed in detail in the next section.